OpenStudy (anonymous):

find the derivative of x^2/sqrt(x-1) i got (x^2 −2x)/(x−1)^(3/2) is this right? someone plz check

OpenStudy (dumbcow):

i get \[\frac{3x^{2}-4x}{2(x-1)^{3/2}}\]

OpenStudy (anonymous):

ugh what am i doing wrong? lol

OpenStudy (dumbcow):

this might help http://www.wolframalpha.com/input/?i=differentiate+x^2%2Fsqrt%28x-1%29

OpenStudy (dumbcow):

steps may be a little confusing though

OpenStudy (anonymous):

i guess what im having touble with is g'(x)

OpenStudy (dumbcow):

\[d/dx \sqrt{x-1} = \frac{1}{2\sqrt{x-1}}\]

OpenStudy (anonymous):

one question

OpenStudy (anonymous):

does the two come from the exponent of x

OpenStudy (dumbcow):

yes, the exponent is 1/2

OpenStudy (anonymous):

ok thank you for the help